# MentalMAC: Enhancing Large Language Models for Detecting Mental Manipulation via Multi-Task Anti-Curriculum Distillation

### v1

@@ Abstract
Mental manipulation is a subtle yet pervasive form of psychological abuse that poses serious threats to mental health. Its covert nature and the complexity of manipulation strategies make it challenging to detect, even for state-of-the-art large language models (LLMs). This concealment also hinders the manual collection of large-scale, high-quality annotations essential for training effective models. Although recent efforts have sought to improve LLM’s performance on this task, progress remains limited due to the scarcity of real-world annotated datasets. To address these challenges, we propose MentalMAC, a multi-task anti-curriculum distillation method that enhances LLMs’ ability to detect mental manipulation in multi-turn dialogue. Our approach includes: (i) EvoSA, an unsupervised data expansion method based on evolutionary operations and speech act theory; (ii) teacher-model-generated multi-task supervision; and (iii) progressive knowledge distillation from complex to simpler tasks. We then constructed the ReaMent dataset with 5,000 real-world dialogue samples, using a MentalMAC-distilled model to assist human annotation. Vast experiments demonstrate that our method significantly narrows the gap between student and teacher models and outperforms competitive LLMs across key evaluation metrics. All code, datasets, and checkpoints will be released upon paper acceptance.
@@ 1 Introduction
Mental manipulation is a subtle and covert form of psychological control that employs language to influence and alter another person’s perception for the manipulator’s own benefit. This type of manipulation is widespread in interpersonal relationships and can have serious negative impacts on mental health (Wang et al., 2024). On social networks, perpetrators often use deception and intimidation to coerce children into engaging in harmful online behaviors, which can lead to long-lasting psychological consequences such as anxiety, depression, and self-harm (Ramiro et al., 2019; World Health Organization, 2024). Within intimate relationships, manipulative tactics like gaslighting systematically distort victims’ sense of reality, resulting in emotional instability and an increased risk of suicidal thoughts (Lohmann et al., 2024; Munro et al., 2024). Furthermore, as AI chatbots become increasingly sophisticated and interactive, new cases of mental manipulation have arisen, with reports highlighting the induction of dependency in adults and the troubling normalization of violent ideologies among children (National Public Radio, 2024; Laestadius et al., 2024).
@@
Unlike other harmful content, mental manipulation is often concealed within a multi-turn dialogue, making it difficult to detect even for experts (Myhill and Hohl, 2019). Wang et al. (2024) introduced the MentalManip dataset and found manipulative dialogues are semantically similar to non-manipulative ones, highlighting the challenge of detection based solely on lexical or semantic cues. Although large language models (LLMs) have advanced many tasks (Yuan et al., 2024), several state-of-the-art (SOTA) LLMs still perform poorly on this task (see Figure 1). First, data resources are limited and difficult to expand due to the covert nature of manipulation; despite LLMs’ strong rewriting abilities, manual annotation remains necessary due to their misalignment with human cognition (Wang et al., 2024). Second, LLMs struggle to understand deep semantics and speech acts in conversations, often confusing manipulative with non-manipulative interactions involving persuasion or arguments. While chain-of-thought (CoT) (Yang et al., 2024) and intent-aware prompting (Ma et al., 2025) provide some improvement, gains remain minimal. Third, the sole dataset for this task, MentalManip, is derived from the Cornell Movie Dialogs Corpus [^1] https://www.cs.cornell.edu/~cristian/Cornell_Movie-Dialogs_Corpus.html and contains only 4,000 samples. The scarcity of real-world dialogue datasets significantly limits progress.
@@
Figure 2: Overall workflow of the proposed MentalMAC. In Step 1, we introduce EvoSA, a novel prompting method designed for the unsupervised generation of dialogues containing or not containing mental manipulation. In Step 2, we employ a teacher model to generate both correct and incorrect rationales, along with feedback for the incorrect ones, to construct multi-task training data. In Step 3, we apply anti-curriculum distillation, training the student model from hard to easy tasks to enhance its performance.
@@
Given the limited data in this domain, we propose ReaMent [^4] Real-world Mental manipulation, a dataset of 5,000 real-world dialogues curated from YTD-18M (Han et al., 2023) for mental manipulation detection. To reduce manual labor, we leveraged a student model distilled via MentalMAC to perform initial screening over a large set of dialogues, identifying those that potentially contain elements of mental manipulation. Human annotators then reviewed and corrected the results to ensure label accuracy. In contrast to MentalManip, which uses scripted movie dialogues, ReaMent offers a more realistic representation of mental manipulation in natural settings. Our key contributions are summarized as follows:
  • MentalMAC enhances LLMs’ ability to detect mental manipulation through three stages: First, EvoSA enables unsupervised data expansion by teachers; Second, the teacher generates multi-task data; Third, tasks are arranged from hard to easy to enable anti-curriculum distillation for student models.
@@ 2.1 Mental Manipulation Detection
Mental manipulation has recently drawn increased attention due to its subtle and insidious nature. However, only one dataset, MentalManip, exists for this task, containing just 4,000 samples. This limited resource highlights the urgent need for larger, high-quality datasets, motivating the construction of ReaMent. Current detection methods mostly rely on large LLMs or commercial APIs with prompting strategies such as few-shot learning and chain-of-thought, which incur high costs (Yang et al., 2024). Although Ma et al. (2025) proposed intent-aware prompting to improve LLMs’ understanding of dialogue intent, improvements remain limited. The scarcity of targeted research reveals a significant gap in this area. To address it, we propose MentalMAC, a method that distills the capabilities of large teacher models into smaller student models, enhancing student performance and potentially enabling it to surpass the teacher.
@@ 2.2 Knowledge Distillation for LLMs
Knowledge distillation (KD) (Hinton et al., 2015) is a common technique for transferring knowledge—from a larger teacher model to a smaller student—using hard or soft labels. With Chain-of-Thought (CoT) prompting enhancing LLM reasoning, studies like PINTO (Wang and Chang, 2022), Distilling Step-by-Step (Hsieh et al., 2023), RevThink (Chen et al., 2024), and SuperCorrect (Yang et al., 2025a) have explored multi-task fine-tuning that leverages reasoning explanations and LLM-generated task content. For toxic text detection, Zhang et al. (2024) proposed BD-LLM, which improves detection via bootstrapped prompting and rationale-based distillation. However, BD-LLM depends solely on predefined toxicity definitions and uses a simple rationale distillation, limiting its adaptation to the covert, context-sensitive nature of mental manipulation detection. In contrast, MentalMAC is tailored for detecting mental manipulation in multi-turn dialogues, requiring the student model to master not only correct reasoning but also the identification of errors in deceptive rationales. Inspired by anti-curriculum learning (Wu et al., 2021), we design a progressive distillation schedule that guides the student model from more complex feedback tasks to simpler binary classification, thereby better aligning the training process with the intricate cognitive demands of this task.
@@ 3.1 EvoSA-Based Data Expansion
To address the challenge of lack of sufficient training data, we propose EvoSA, inspired by evolutionary operations (Yang et al., 2025b) and speech act theory (Searle et al., 1980), to drive LLMs to unsupervised data augmentation.
@@
Evolutionary algorithms rely on selection, crossover, and mutation to maintain diversity and adaptability within populations of generated individuals. Besides, speech act theory reveals how the intentions behind our words shape communication in varied contexts, reflecting the rich diversity of conversational goals. Drawing inspiration from both, we introduce EvoSA (Figure 3). The process begins by randomly sampling training dialogues that share the same label, which we call parent dialogues. Guided by evolutionary operations and grounded in speech act theory, LLMs then create novel, distinct child dialogues through mutation and recombination (steps 1–3). To guarantee the quality and consistency of labels, the LLM is tasked with polishing the child dialogue, analyzing the reasoning behind the parent label, and revising the child dialogue accordingly to ensure it remains aligned with the original intent (steps 4–7).
@@
From Equation 2, EvoSA can expand nearly 5,000 high-quality dialogues from just 100 labeled ones. To further validate the generated data, we manually evaluated a random sample of the generated dialogues. Examples and evaluation results are provided in Appendix A.
@@ 3.2 Multi-Task Data Generation
where $P_{ir}$ is a prompt template used to drive the teacher model to generate the incorrect rationale. $r_{i}^{-}$ is the incorrect rationale corresponding to the $i$-th sample $(x_{i}^{\prime},y_{i}^{\prime})$. Afterwards, we prompt the teacher model to generate feedback for incorrect answers, denoted as
@@
where $P_{f}$ is a prompt template used to drive the teacher model to generate feedback on an incorrect rationale. $f_{i}$ is the feedback corresponding to the $i$-th incorrect rationale.
@@
For Task 3, only the label is required. Prompt templates used for multi-task data generation are detailed in Appendix B.
@@ 3.3 Anti-Curriculum Distillation for Detecting Mental Manipulation
Anti-curriculum learning, which models human learning from hard to easy tasks, has proven effective despite seeming counterintuitive (Wu et al., 2021; Liu et al., 2022; Jarca et al., 2025). However, simultaneously learning multiple tasks with varying difficulty levels often leads to imbalanced training progress, which can negatively impact the performance of the primary task. To address this challenge and enhance the model’s effectiveness in detecting mental manipulation in multi-turn dialogues, a hard-to-easy distillation strategy, namely anti-curriculum distillation, is adopted as illustrated in Figure 2.
In the first stage, we prompt the student model $\mathcal{S}$ to act as a teacher by providing it with incorrect rationales without explicitly indicating the error, and asking it to identify the flaw and provide feedback. Mathematically, this can be formulated as
@@
After the first stage, the student model gains an initial understanding of mental manipulation detection. In the second stage, it is further trained on both Task 1 and Task 2. For Task 2, the model must judge if a dialogue contains mental manipulation and provide a rationale, as formalized below:
@@
Overall, Task 1 is learned across all three stages, Task 2 across two stages, and Task 3 only in the final stage. This design ensures that more cognitively demanding tasks receive greater training exposure, thereby enhancing the model’s overall capability. By first engaging in Task 1 and Task 2, the student model learns to identify flaws in incorrect rationales and articulate correct reasoning. This foundation enables more effective learning of Task 3, which requires concise yet accurate judgments without explanatory support.
@@ 5 Experiments
Table 1: Performance of large LLMs and student models on MentalManip and ReaMent. Acc = Accuracy; Pre = Precision; Re = Recall; F1mac/w = macro- and weighted-F1. * indicates data from original papers. "Claude-3.5-Sonnet" and "Qwen2.5-3B" refer to Claude-3.5-Sonnet-20241022 and Qwen2.5-3B-Instruct, respectively. "IAP" and "Distill" stand for intent-aware prompting and distilling step-by-step. Llama3-70B-Instruct is the sole teacher model used. The best and second-best student model results are bolded and underlined, respectively.
@@ 5.1 Experiment Setting
Implementation Details. All experiments were conducted on two NVIDIA H800 GPUs using the Lion optimizer (Chen et al., 2023) and LoRA (Hu et al., 2022). Training was performed for 6 epochs on MentalManip and 5 epochs on ReaMent, with LoRA configured at rank 8, a learning rate of 5e-6, batch size of 8, and gradient accumulation set to 4. During inference, we used a temperature of 0.1 and top-p of 0.5. Additional implementation details can be found in Appendix D.
@@ 5.4 Cross-Dataset Generalization Analysis
Interestingly, we observe that the model trained on ReaMent achieves better results on MentalManip than on ReaMent itself. This may be attributed to the nature of MentalManip’s movie-based dialogues, which tend to be more dramatized yet simpler, while ReaMent contains richer and more complex real-world conversations. Consequently, training on ReaMent enables the model to capture more generalizable features.
@@ 6 Conclusion
We propose MentalMAC, a novel method that enhances LLMs’s ability to detect mental manipulation in multi-turn dialogues via multi-task anti-curriculum distillation. To overcome the challenge of limited training data, we introduce EvoSA, an unsupervised data expansion strategy grounded in evolutionary operations and speech act theory. Our approach defines three progressively simpler tasks: providing feedback on incorrect rationales, generating correct answers with rationales, and producing answers only. A teacher model generates training samples for each task, while the student model is trained in a hard-to-easy order. Leveraging the distilled student model, we support human annotation efforts and construct ReaMent, a dataset of 5,000 real-world dialogues that addresses a critical gap in mental manipulation research. Future work will explore more efficient distillation strategies to further enhance model performance.
@@
EvoSA Our analysis of the data synthesized by EvoSA (see Appendix A for details) reveals that the quality of generated samples may degrade when the parent dialogue is too short, indicating a dependency on sufficient contextual input. Furthermore, we did not explore how EvoSA performs across different LLMs, leaving open the question of model-specific effectiveness. Lastly, due to the safety constraints imposed by LLMs, EvoSA is currently unable to generate dialogues that simultaneously contain toxic content and elements of mental manipulation.
ReaMent Real-world dialogues are often multi-party, whereas our ReaMent dataset contains only two-person conversations. Although we found that EvoSA can extend two-person dialogues into multi-party ones (see Appendix A for details), this remains a limitation of the ReaMent dataset. Additionally, ReaMent is currently limited to English content, which may restrict its applicability in multilingual or cross-cultural contexts.

### v2

@@ Abstract
Mental manipulation is a subtle yet pervasive form of psychological abuse that poses serious threats to mental health. Its covert nature and the complexity of manipulation strategies make it challenging to detect, even for state-of-the-art large language models (LLMs). This concealment also hinders the manual collection of large-scale, high-quality annotations essential for training effective models. Although recent efforts have sought to improve LLMs’ performance on this task, progress remains limited due to the scarcity of real-world annotated datasets. To address these challenges, we propose MentalMAC, a multi-task anti-curriculum distillation method that enhances LLMs’ ability to detect mental manipulation in multi-turn dialogue. Our approach includes: (i) EvoSA, an unsupervised data expansion method based on evolutionary operations and speech act theory; (ii) teacher model-generated multi-task supervision; and (iii) progressive knowledge distillation from complex to simpler tasks. We then constructed the ReaMent dataset with 5,000 real-world dialogue samples, using a MentalMAC-distilled model to assist human annotation. Vast experiments demonstrate that our method significantly narrows the gap between student and teacher models and outperforms competitive LLMs across key evaluation metrics. All code, datasets, and checkpoints will be released upon paper acceptance.
@@ 1 Introduction
Mental manipulation is a subtle and covert form of psychological control that employs language to influence and alter another person’s perception for the manipulator’s benefit. This type of manipulation is widespread in interpersonal relationships and can have severe negative impacts on mental health (Wang et al., 2024). On social networks, perpetrators often use deception and intimidation to coerce children into engaging in harmful online behaviors, which can lead to long-lasting psychological consequences such as anxiety, depression, and self-harm (Ramiro et al., 2019; World Health Organization, 2024). Within intimate relationships, manipulative tactics like gaslighting systematically distort victims’ sense of reality, resulting in emotional instability and an increased risk of suicidal thoughts (Lohmann et al., 2024; Munro et al., 2024). Furthermore, as AI chatbots become increasingly sophisticated and interactive, new cases of mental manipulation have arisen, with reports highlighting the induction of dependency in adults and the troubling normalization of violent ideologies among children (National Public Radio, 2024; Laestadius et al., 2024).
@@
Unlike other harmful content, mental manipulation is often concealed within a multi-turn dialogue, making it difficult to detect even for experts (Myhill and Hohl, 2019). Wang et al. (2024) introduced the MentalManip dataset and found that manipulative dialogues are semantically similar to non-manipulative ones, highlighting the challenge of detection based solely on lexical or semantic cues. Although large language models (LLMs) have advanced many tasks (Yuan et al., 2024), several state-of-the-art (SOTA) LLMs still perform poorly on this task (see Figure 1). First, data resources are limited and difficult to expand due to the covert nature of manipulation; despite LLMs’ strong rewriting abilities, manual annotation remains necessary due to their misalignment with human cognition (Wang et al., 2024). Second, LLMs struggle to understand deep semantics and speech acts in conversations, often confusing manipulative with non-manipulative interactions involving persuasion or arguments. While chain-of-thought (CoT) (Yang et al., 2024) and intent-aware prompting (Ma et al., 2025) provide some improvement, gains remain minimal. Third, the sole dataset for this task, MentalManip, is derived from the Cornell Movie Dialogs Corpus [^1] https://www.cs.cornell.edu/~cristian/Cornell_Movie-Dialogs_Corpus.html and contains only 4,000 samples. The scarcity of real-world dialogue datasets significantly limits progress.
@@
Figure 2: Overall workflow of the proposed MentalMAC. In Step 1, we introduce EvoSA, a novel prompting method designed for the unsupervised generation of dialogues containing or not containing mental manipulation. In Step 2, we employ a teacher model to generate correct and incorrect rationales and feedback for incorrect ones to construct multi-task training data. In Step 3, we apply anti-curriculum distillation, training the student model from hard to easy tasks to enhance its performance.
@@
Given the limited data in this domain, we propose ReaMent [^4] Real-world Mental manipulation, a dataset of 5,000 real-world dialogues curated from YTD-18M (Han et al., 2023) for mental manipulation detection. To reduce manual labor, we leveraged a student model distilled via MentalMAC to perform initial screening over a large set of dialogues, identifying those potentially containing mental manipulation elements. Human annotators then reviewed and corrected the results to ensure label accuracy. In contrast to MentalManip, which uses scripted movie dialogues, ReaMent offers a more realistic representation of mental manipulation in natural settings. Our key contributions are summarized as follows:
  • MentalMAC enhances LLMs’ ability to detect mental manipulation through three stages: First, EvoSA enables unsupervised data expansion by teachers; Second, the teacher generates multi-task data; Third, tasks are arranged from hard to easy to allow anti-curriculum distillation for student models.
@@ 2.1 Mental Manipulation Detection
Mental manipulation has recently drawn increased attention due to its subtle and insidious nature. However, only one dataset, MentalManip, exists for this task, containing just 4,000 samples. This limited resource highlights the urgent need for larger, high-quality datasets, motivating the construction of ReaMent. Current detection methods mainly rely on large LLMs or commercial APIs with prompting strategies such as few-shot learning and chain-of-thought, which incur high costs (Yang et al., 2024). Although Ma et al. (2025) proposed intent-aware prompting to improve LLMs’ understanding of dialogue intent, improvements remain limited. The scarcity of targeted research reveals a significant gap in this area. To address it, we propose MentalMAC, a method that distills the capabilities of large teacher models into smaller student models, enhancing student performance and potentially enabling it to surpass the teacher.
@@ 2.2 Knowledge Distillation for LLMs
Knowledge distillation (KD) (Hinton et al., 2015) is a common technique for transferring knowledge from a larger teacher model to a smaller student, using hard or soft labels. With Chain-of-Thought (CoT) prompting enhancing LLM reasoning, studies like PINTO (Wang and Chang, 2022), Distilling Step-by-Step (Hsieh et al., 2023), RevThink (Chen et al., 2024), and SuperCorrect (Yang et al., 2025a) have explored multi-task fine-tuning that leverages reasoning explanations and LLM-generated task content. For toxic text detection, Zhang et al. (2024) proposed BD-LLM, which improves detection via bootstrapped prompting and rationale-based distillation. However, BD-LLM depends solely on predefined toxicity definitions and uses a simple rationale distillation, limiting its adaptation to the covert, context-sensitive nature of mental manipulation detection. In contrast, MentalMAC is tailored for detecting mental manipulation in multi-turn dialogues, requiring the student model to master correct reasoning and the identification of errors in deceptive rationales. Inspired by anti-curriculum learning (Wu et al., 2021), we design a progressive distillation schedule that guides the student model from more complex feedback tasks to simpler binary classification, thereby better aligning the training process with the intricate cognitive demands of this task.
@@ 3.1 EvoSA-Based Data Expansion
To address the challenge of a lack of sufficient training data, we propose EvoSA, inspired by evolutionary operations (Yang et al., 2025b) and speech act theory (Searle et al., 1980), to drive LLMs to unsupervised data augmentation.
@@
Evolutionary algorithms rely on selection, crossover, and mutation to maintain diversity and adaptability within populations of generated individuals. Besides, speech act theory reveals how the intentions behind our words shape communication in varied contexts, reflecting the rich diversity of conversational goals. Drawing inspiration from both, we introduce EvoSA (Figure 3). The process begins by randomly sampling training dialogues with the same label, which we call parent dialogues. Guided by evolutionary operations and grounded in speech act theory, LLMs then create novel, distinct child dialogues through mutation and recombination (steps 1–3). To guarantee the quality and consistency of labels, the LLM is tasked with polishing the child dialogue, analyzing the reasoning behind the parent label, and revising the child dialogue accordingly to ensure it remains aligned with the original intent (steps 4–7).
@@
From Equation 2, EvoSA can expand nearly 5,000 high-quality dialogues from just 100 labeled ones. We manually evaluated a random sample of the generated dialogues to validate the generated data further. Examples and evaluation results are provided in Appendix A.
@@ 3.2 Multi-Task Data Generation
where $P_{ir}$ is a prompt template that drives the teacher model to generate the incorrect rationale. $r_{i}^{-}$ is the incorrect rationale corresponding to the $i$-th sample $(x_{i}^{\prime},y_{i}^{\prime})$. Afterwards, we prompt the teacher model to generate feedback for incorrect answers, denoted as
@@
where $P_{f}$ is a prompt template that drives the teacher model to generate feedback on an incorrect rationale. $f_{i}$ is the feedback corresponding to the $i$-th incorrect rationale.
@@
For Task 3, only the label is required. Prompt templates for multi-task data generation are detailed in Appendix B.
@@ 3.3 Anti-Curriculum Distillation for Detecting Mental Manipulation
Anti-curriculum learning models human learning from hard to easy tasks, has proven effective despite seeming counterintuitive (Wu et al., 2021; Liu et al., 2022; Jarca et al., 2025). However, simultaneously learning multiple tasks with varying difficulty levels often leads to imbalanced training progress, which can negatively impact the performance of the primary task. To address this challenge and enhance the model’s effectiveness in detecting mental manipulation in multi-turn dialogues, a hard-to-easy distillation strategy, namely anti-curriculum distillation, is adopted as illustrated in Figure 2.
In the first stage, we prompt the student model $\mathcal{S}$ to act as a teacher by providing incorrect rationales without explicitly indicating the error, and asking it to identify the flaw and provide feedback. Mathematically, this can be formulated as
@@
After the first stage, the student model initially understands mental manipulation detection. In the second stage, it is further trained on Task 1 and Task 2. For Task 2, the model must judge if a dialogue contains mental manipulation and provide a rationale, as formalized below:
@@
Overall, Task 1 is learned across all three stages, Task 2 across two, and Task 3 only in the final stage. This design ensures that more cognitively demanding tasks receive greater training exposure, enhancing the model’s overall capability. By first engaging in Tasks 1 and 2, the student model learns to identify flaws in incorrect rationales and articulate correct reasoning. This foundation enables more effective learning of Task 3, which requires concise yet accurate judgments without explanatory support.
@@ 5 Experiments
Table 1: Performance of large LLMs and student models on MentalManip and ReaMent. Acc = Accuracy; Pre = Precision; Re = Recall; F1mac/w = macro- and weighted-F1. * indicates data from original papers. "Claude-3.5-Sonnet" and "Qwen2.5-3B" refer to Claude-3.5-Sonnet-20241022 and Qwen2.5-3B-Instruct, respectively. "IAP" and "Distill" mean intent-aware prompting and distilling step-by-step. Llama3-70B-Instruct is the sole teacher model used. The best and second-best student model results are bolded and underlined, respectively.
@@ 5.1 Experiment Setting
Implementation Details. All experiments were conducted on two NVIDIA H800 GPUs using the Lion optimizer (Chen et al., 2023) and LoRA (Hu et al., 2022). Training was performed for six epochs on MentalManip and five epochs on ReaMent, with LoRA configured at rank 8, a learning rate of 5e-6, batch size of 8, and gradient accumulation set to 4. During inference, we used a temperature of 0.1 and top-p of 0.5. Additional implementation details can be found in Appendix D.
@@ 5.4 Cross-Dataset Generalization Analysis
Interestingly, we observe that the model trained on ReaMent achieves better results on MentalManip than on ReaMent itself. This may be attributed to the nature of MentalManip’s movie-based dialogues, which are more dramatized yet simpler. At the same time, ReaMent contains richer and more complex real-world conversations. Consequently, training on ReaMent enables the model to capture more generalizable features.
@@ 6 Conclusion
We propose MentalMAC, a novel method that enhances LLMs’ ability to detect mental manipulation in multi-turn dialogue via multi-task anti-curriculum distillation. To overcome the challenge of limited training data, we introduce EvoSA, an unsupervised data expansion strategy grounded in evolutionary operations and speech act theory. Our approach defines three progressively simpler tasks: providing feedback on incorrect rationales, generating correct answers with rationales, and producing binary judgements only. A teacher model generates training samples for each task, while the student model is trained in a hard-to-easy order. Leveraging the MentalMAC-distilled student model, we support human annotation efforts and construct ReaMent, a dataset of 5,000 real-world dialogues that addresses a critical gap in mental manipulation research. Future work will explore more efficient distillation strategies to enhance model performance further in this task.
@@
EvoSA Our analysis of the data synthesized by EvoSA (see Appendix A for details) reveals that the quality of generated samples may degrade when the parent dialogue is too short, indicating a dependency on sufficient contextual input. Furthermore, we did not explore how EvoSA performs across different LLMs, leaving open the question of model-specific effectiveness. Lastly, due to the safety constraints imposed by LLMs, EvoSA cannot generate dialogues that simultaneously contain toxic content and elements of mental manipulation.
ReaMent Real-world dialogues are often multi-party, whereas our ReaMent dataset contains only two-person conversations. Although we found that EvoSA can extend two-person dialogues into multi-party ones (see Appendix A for details), this remains a limitation of the ReaMent dataset. ReaMent is also limited to English content, which may restrict its applicability in multilingual or cross-cultural contexts.